OpenStudy (openstudyuser):

I need help with this concept of Continuity and Discontinuity. Determine whether each function is continuous at the given x value. Justify your answer using the continuity test. y=x^3-4; x=1

OpenStudy (openstudyuser):

Also i would gladly appreciate it if someone could explain the Continuity Test

OpenStudy (turingtest):

\(f(x)\) is continuous at \(x=a\) if and only if\[~~~~~~~~~\lim_{x\to a}f(x)=f(a)\]

OpenStudy (turingtest):

that is, is the limit at that point the same as the value of the function at that point?

OpenStudy (turingtest):

what is\[\lim_{x\to1}x^2-4\]?

OpenStudy (openstudyuser):

Um..im sorry, but i havent learned lim. Would you mind explaining what that means?

OpenStudy (turingtest):

oh, that is kinda hard. I'm not sure what kind of continuity test you are referring to then.

OpenStudy (turingtest):

what is the definition you have?

OpenStudy (openstudyuser):

Iam in precalculus. I just started the course.

OpenStudy (turingtest):

well the limit is the value that the function approaches as x gets close to the specified number, in this case as x approaches 1

OpenStudy (openstudyuser):

We have infinite disconituty, Jump Discontinuity, Point Discontinuity, and Everywhere Discontious.

OpenStudy (turingtest):

yeah you need limits for all that, and I can't explain all the subtleties of the concept that quickly. I suggest you read something like this: http://tutorial.math.lamar.edu/Classes/CalcI/LimitsIntro.aspx and come back with any questions you may have after.

OpenStudy (openstudyuser):

Im looking at my textbook right now, and we have nothing about limits. Thats whats confusing me.

OpenStudy (turingtest):

well if you want to be lazy about it you can simply say "is f(x) ever divided by zero, does it ever have a square root of a negative number, or the logarithm of a negative number or zero?" if the answer is no, and it is not a piecewise function, then f(x) is continuous everywhere

OpenStudy (openstudyuser):

Hahaha, im not lazy about it. I was just confused. Our teacher didnt explain it well.

OpenStudy (turingtest):

No I was just kidding, didn't mean to call you lazy. It's just an incomplete approach.

OpenStudy (openstudyuser):

Okay..I will research on that topic a little later. How would you describe the end behavior of a function then? Like if you had y=x^3 +2x^2+x-1

OpenStudy (openstudyuser):

Does that involve limits too?

OpenStudy (turingtest):

yes, but this is a little easier to ignore the limit bit in the ends the largest power of x takes over, in this case x^3 will win out

OpenStudy (turingtest):

so as x gets large and positive, x^3 gets large and positive, going off to infinity

OpenStudy (turingtest):

in the notation of limits we would write\[\lim_{x\to\infty}x^3+2x^2+x-1=\infty\]that is "the limit as x goes to infinity of f(x) is (positive) infinity

OpenStudy (turingtest):

what about when x is large and negative?

OpenStudy (turingtest):

\[\lim_{x\to-\infty}x^3+2x^2+x-1=?\]

OpenStudy (openstudyuser):

Im gonna find out how to use limits. Because im not understanding this at all. I appreciate your help. I will find out how to do it, and come back to this

OpenStudy (turingtest):

okay, good luck